# Lost in the FoG: Pitfalls of Models for Large-Scale Hydrogen Distributions

### v4

@@ Abstract
Large-scale ${\rm H}$i surveys and their cross-correlations with galaxy distributions have immense potential as cosmological probes. Interpreting these measurements requires theoretical models that must incorporate redshift-space distortions (RSDs), such as the Kaiser and fingers-of-God (FoG) effect, and differences in the tracer and matter distributions via the tracer bias. These effects are commonly approximated with assumptions that should be tested on simulated distributions. In this work, we use the hydrodynamical simulation suite IllustrisTNG to assess the performance of models of $z\leq 1$ ${\rm H}$i auto and ${\rm H}$i-galaxy cross-power spectra, finding that the models employed by recent observations introduce errors comparable to or exceeding their measurement uncertainties. In particular, neglecting FoG causes $\gtrsim 10\%$ deviations between the modeled and simulated power spectra at $k\gtrsim 0.1$ h cMpc-1, larger than assuming a constant bias which reaches the same error threshold at slightly smaller scales. However, even without these assumptions, models can still err by $\sim 10\%$ on relevant scales. These remaining errors arise from multiple RSD damping sources on ${\rm H}$i clustering, which are not sufficiently described with a single FoG term. Overall, our results highlight the need for an improved understanding of RSDs to harness the capabilities of future measurements of ${\rm H}$i distributions.
@@ 1 Introduction
Cosmology imprints on the large-scale structure of the Universe, allowing measurements of the distribution of matter to probe the nature of dark matter and dark energy and constrain cosmological parameters (Jenkins et al., 1998; Eisenstein et al., 2005; Reddick et al., 2014; DESI Collaboration et al., 2025). However, we cannot observe the Universe’s large-scale structure directly. We instead rely on surveys of baryonic tracers such as atomic neutral hydrogen (${\rm H}$i) or galaxies to infer the underlying matter distribution (DESI Collaboration et al., 2024), a process which requires models of the matter-tracer relationship. These models are simple in the linear regime where tracers faithfully follow the matter distribution, but baryonic physics and nonlinear effects complicate the behavior on small scales, demanding more sophisticated models. Despite the challenge, the stronger constraining power on small scales makes the development of these complex models worthwhile (Krause & Eifler, 2017; Chisari et al., 2019).
@@
Consequently, a large body of work has been dedicated to improving our understanding of structure in the quasi-linear and nonlinear regime. Notable examples of modeling techniques include Lagrangian perturbation theory (LPT, Bernardeau et al., 2002; Carlson et al., 2013; Vlah et al., 2015) and halo occupation distribution (HOD) models (Zehavi et al., 2004; Zheng et al., 2007). However, both techniques have limitations: LPT does not natively include baryonic effects and HOD requires knowledge and assumptions of how tracers occupy halos. In certain applications, a simpler and more general prescription which assumes parameterized forms for the various nonlinear contributions is preferred.
@@
These general prescriptions have interpreted measurements of ${\rm H}$i and galactic distributions to constrain cosmological parameters. Of the two tracers, galaxies are the traditional probe of large-scale structure (e.g., Cole et al., 2005), but ${\rm H}$i has arisen as a promising candidate due to its ability to cheaply and quickly observe large volumes of space via 21cm intensity mapping (Bharadwaj et al., 2001; Battye et al., 2004; Leo et al., 2020). However, 21cm intensity mapping experiments still struggle with low signal-to-noise ratios, arising from galactic foregrounds and systematics (Matteo et al., 2002), making the interpretation of ${\rm H}$i auto power spectra challenging (Paul et al., 2023, hereafter P23). These noise sources are absent in galaxy surveys, so the noise can be reduced by cross-correlating ${\rm H}$i and galaxies. These cross-correlations have constrained the cosmological abundance of ${\rm H}$i ($\Omega_{\mathrm{HI}}$) at low redshifts ($z\leq 1$, Chang et al., 2010; Masui et al., 2013; Switzer et al., 2013; Anderson et al., 2018). In particular, Wolz et al. (2022, hereafter W22), Cunnington et al. (2022, C22), CHIME Collaboration et al. (2023, CHIME23), and Carucci et al. (2024, C24) have produced $\Omega_{\mathrm{HI}}$ constraints competitive to other ${\rm H}$i observation techniques. As measurement and data processing techniques improve (Cunnington et al., 2019; Carucci et al., 2020; Podczerwinski & Timbie, 2024), ${\rm H}$i-galaxy cross-correlations will probe other cosmological quantities in the near future (MeerKLASS Collaboration et al., 2024).
@@
However, the resulting constraints depend on how well the adopted models capture the relationship between observed tracer and matter distributions. We can evaluate these models’ performance by comparing their predictions to known ${\rm H}$i, galaxy, and matter distributions in simulations. Cosmological hydrodynamical simulations now produce realistic low-redshift ${\rm H}$i and galaxy distributions in sufficiently large volumes to be useful for this purpose (Bahé et al., 2016; Nelson et al., 2018; Diemer et al., 2019; Stevens et al., 2019; Davé et al., 2020). In particular, Osinga et al. (2024) showed that the ${\rm H}$i auto- and cross-correlations from IllustrisTNG agree closely with relevant observations at $z\leq 1$. We leverage the agreement between IllustrisTNG and observations to gain insight into the performance of large-scale ${\rm H}$i models in two phases. First, we report IllustrisTNG’s predictions for the various model components and assess how well their simplified forms in the general model agree with the actual values. Second, we evaluate the accuracy of popular large-scale prescriptions by comparing their predictions to the correlations computed directly from IllustrisTNG.
@@ 2 Simulation data
We use the IllustrisTNG suite of cosmological magneto-hydrodynamics simulations (Nelson et al., 2018; Pillepich et al., 2018a; Springel et al., 2018; Naiman et al., 2018; Marinacci et al., 2018; Nelson et al., 2019). This suite offers simulations in three different box sizes, 35 h-1 cMpc, 75 h-1 cMpc, and 205 h-1 cMpc, at varying resolutions, using the Planck Collaboration et al. (2016) cosmological parameters ($\Omega_{\rm{m}}=0.3089$, $\Omega_{\rm{b}}=0.0486$, $h=0.6774$, $\sigma_{\rm{s}}=0.8159$).
@@
The simulation is evolved using the AREPO code (Springel, 2010), which employs a tree-PM method for gravity calculations and a Voronoi mesh-based Godunov scheme for magneto-hydrodynamics. The IllustrisTNG simulations incorporate sub-grid models to simulate unresolved processes like star formation, gas cooling, and AGN (Vogelsberger et al., 2013; Weinberger et al., 2018). The models are calibrated against a selection of observational data to accurately represent the galaxy population at low redshifts (Pillepich et al., 2018b). Dark matter haloes are identified using the Friends-of-Friends (FoF) algorithm (Davis et al., 1985), and their internal substructures, or “subhalos”, are cataloged using the Subfind algorithm (Springel et al., 2001).
@@
Our main analyses focus on the largest box, 205 h-1 cMpc (TNG300), at redshifts $z=0,0.5,$ and $z=1$. We discuss the impact of resolution in Section 6 and directly compare our results to the 75 h-1 cMpc box (TNG100) in Appendix A. To limit the impact of poorly resolved galaxies, we restrict our galaxy population to those containing at least 200 stellar particles, corresponding to a minimum stellar mass of $\approx 10^{9}M_{\odot}$. We limit our analysis to redshifts $z=0,0.5,$ and $1$ where the color distribution in IllustrisTNG matches observations. The agreement worsens at earlier redshifts because IllustrisTNG lacks red, star-forming, dusty galaxies (Donnari et al., 2019; Gebek et al., 2023, 2025).
@@
We separate our galaxy sample into blue and red populations to mimic how galaxies are selected in optical surveys, as they detect galaxy samples visible in certain optical bands such as blue emission-line galaxies (ELGs) and luminous red galaxies (LRGs) (Alam et al., 2021). We approximate these populations by categorizing galaxies as red or blue with their rest-frame $g\!-\!r$ color values (Stoughton et al., 2002): galaxies with $g\!-\!r\geq 0.6$ are blue, with all others being red. This threshold changes with time to reflect the overall color evolution of the galaxy population. Specifically, $g\!-\!r$ thresholds are set at 0.6, 0.55, and 0.5 for redshifts 0, 0.5, and 1, respectively. While testing various color definitions with and without dust corrections (Nelson et al., 2018), Osinga et al. (2024) found that these adjustments only substantially impact small-scale clustering at $k\gtrsim 1$ h cMpc-1 (see their appendix A). A more exact method of matching IllustrisTNG galaxies to ELGs and LRGs will only somewhat impact the quantitative results from Section 4 (e.g., Tables 1-3). The scale-dependent behavior in Section 4 and overall model performance presented in Section 5 are insensitive to small changes in the definitions of blue and red galaxies.
@@
The final tracer, ${\rm H}$i, is not explicitly modeled in IllustrisTNG, so we must turn to post-processing to separate atomic and molecular hydrogen. We have tested a wide range models, but find negligible difference between the different treatments (an offset of $\lesssim 2\%$ in amplitude, see Osinga et al., 2024), consequently we restrict our analysis to the single model from Villaescusa-Navarro et al. (2018), hereafter VN18.
@@ 3 Models of large-scale ${\rm H}$i distributions
$\langle\tilde{\delta}_{i}(\boldsymbol{k})\tilde{\delta}_{j}(\boldsymbol{k}^{\prime})\rangle=(2\pi)^{3}P_{i\times j}(k)\delta_{D}^{3}(\boldsymbol{k}-\boldsymbol{k}^{\prime})\,.$ (1)
@@ 3.1 Nonlinearities
At large scales, prescriptions are relatively simple since the growth of the small large-scale perturbations ($\delta\ll 1$) can be described linearly (for derivation, see Dodelson & Schmidt, 2020). However, this growth becomes increasingly nonlinear toward small scales until gravitational instabilities form collapsed structures. On these scales, we rely on numerical methods to predict the matter power spectrum.
The transition between the linear and nonlinear regime can be found by comparing their power spectra in Fig. 1. We use camb (Lewis et al., 2000; Lewis & Challinor, 2011) and halofit (Takahashi et al., 2012; Smith & Angulo, 2019) to determine the linear ($P_{\mathrm{L}}$) and nonlinear ($P_{\mathrm{NL}}$) power spectra, respectively. The two diverge at $k\sim 0.1$ h cMpc-1 at $z=1$, which agrees with other works studying the linear-nonlinear transition (e.g., Smith et al., 2003). Roughly at the same scales, the wiggles from baryonic acoustic oscillations (Eisenstein et al., 2005) imprint on both the linear and nonlinear power spectra.
We also compare the nonlinear power spectra calculated from TNG300 and halofit. At large scales, the two roughly agree, although the small number of modes at the largest scales that TNG300 probes causes statistical variations that fall above and below the halofit result (we address this in Section 4.1). At smaller scales, halofit and TNG300 diverge, with TNG300 having slightly more power, although the shapes match closely. Differences are expected since the halofit results are tuned to dark matter-only N-body simulations and the TNG300 power spectra include baryonic effects and are subject to cosmic variance (Miller et al., 2002; Somerville et al., 2004; van Daalen et al., 2011; Li et al., 2014; Springel et al., 2018). For demonstrative purposes, we will use the halofit matter power spectrum for this section but strictly use the TNG300 result for the remainder of the paper.
@@ 3.2 From matter to tracers
The impact of the ${\rm H}$i and galaxy biases is shown with the purple line in Fig. 1, which corresponds mathematically to $b_{\mathrm{HI}}(k)b_{\mathrm{Gal}}(k)P_{\mathrm{NL}}(k)$. The ratios in the bottom panel of Fig. 1 show that both biases are constant on large scales until $k\sim 0.1$ h cMpc-1 where they increase. The galaxy and ${\rm H}$i biases inherit this shape from the halo bias — halos tend to occupy the most nonlinear regions of space and thus are increasingly more clustered than matter on small scales, such that the halo bias increases (Jeong & Komatsu, 2009; Nishimichi, 2012; Nishizawa et al., 2013; Paranjape et al., 2013). Nearly all galaxies and ${\rm H}$i inhabit halos at $z\leq 1$ (White & Rees, 1978; Villaescusa-Navarro et al., 2018) such that their biases mirror the scale-dependence of the halo bias. The ${\rm H}$i and galaxy biases used in Fig. 1 are estimated from TNG300 (discussed more in Section 4) — for scales beyond what TNG300 probes, we assume that the biases are constant and extrapolate the large-scale values from TNG300.
@@
Similarly to the bias, the correlation coefficient is expected to be constant on large scales and approach zero toward small scales as stochasticity becomes more significant (Carucci et al., 2017; Osinga et al., 2024). These trends are shown with the maroon line in Fig. 1, which represents the expression $b_{\mathrm{HI}}(k)b_{\mathrm{Gal}}(k)r_{\mathrm{HI-Gal}}(k)P_{\mathrm{NL}}(k)$ and completely describes the tracer-tracer and tracer-matter relationships. We use the ${\rm H}$i-galaxy correlation coefficient from TNG300, assuming that $r_{\mathrm{HI-Gal}}=1$ at scales larger than what TNG300 probes. The correlation coefficient suppresses the power spectrum at $k\gtrsim 0.2$ h cMpc-1, with this effect strengthening toward small scales. Overall, the contribution of the correlation coefficient is fairly small relative to the other components, reducing the power spectrum by a factor of 2 at the smallest scales shown in Fig. 1.

### v5

@@ Abstract
Large-scale $\rm H$}i surveys and their cross-correlations with galaxy distributions have immense potential as cosmological probes. Interpreting these measurements requires theoretical models that must incorporate redshift-space distortions (RSDs), such as the Kaiser and fingers-of-God (FoG) effect, and differences in the tracer and matter distributions via the tracer bias. These effects are commonly approximated with assumptions that should be tested on simulated distributions. In this work, we use the hydrodynamical simulation suite IllustrisTNG to assess the performance of models of $z\leq 1$ $\rm H$}i auto and $\rm H$}i-galaxy cross-power spectra, finding that the models employed by recent observations introduce errors comparable to or exceeding their measurement uncertainties. In particular, neglecting FoG causes $\gtrsim 10\%$ deviations between the modeled and simulated power spectra at $k\gtrsim 0.1$ h cMpc-1, larger than assuming a constant bias which reaches the same error threshold at slightly smaller scales. However, even without these assumptions, models can still err by $\sim 10\%$ on observationally relevant scales. These remaining errors arise from multiple RSD damping sources on $\rm H$}i clustering, which are not sufficiently described with a single FoG term. Overall, our results highlight the need for an improved understanding of RSDs to harness the capabilities of future measurements of $\rm H$}i distributions.
[…]
software: numpy (C. R. Harris et al., 2020), matplotlib (J. D. Hunter, 2007), seaborn (M. Waskom, 2021), h5py (A. Collette, 2013), colossus (B. Diemer, 2018) and pyfftw (H. Gomersall, 2021)
@@ 1 Introduction
Cosmology imprints on the large-scale structure of the Universe, allowing measurements of the distribution of matter to probe the nature of dark matter and dark energy and constrain cosmological parameters (A. Jenkins et al., 1998; D. J. Eisenstein et al., 2005; R. M. Reddick et al., 2014; DESI Collaboration et al., 2025). However, we cannot observe the Universe’s large-scale structure directly. We instead rely on surveys of baryonic tracers such as atomic neutral hydrogen ($\rm H$}i) or galaxies to infer the underlying matter distribution (DESI Collaboration et al., 2024), a process which requires models of the matter-tracer relationship. These models are simple in the linear regime where tracers faithfully follow the matter distribution, but baryonic physics and nonlinear effects complicate the behavior on small scales, demanding more sophisticated models. Despite the challenge, the stronger constraining power on small scales makes the development of these complex models worthwhile (E. Krause & T. Eifler, 2017; N. E. Chisari et al., 2019).
@@
Consequently, a large body of work has been dedicated to improving our understanding of structure in the quasi-linear and nonlinear regime. Notable examples of modeling techniques include Lagrangian perturbation theory (LPT, F. Bernardeau et al., 2002; J. Carlson et al., 2013; Z. Vlah et al., 2015) and halo occupation distribution (HOD) models (I. Zehavi et al., 2004; Z. Zheng et al., 2007). However, both techniques have limitations: LPT does not natively include baryonic effects and HOD requires knowledge and assumptions of how tracers occupy halos. In certain applications, a simpler and more general prescription which assumes parameterized forms for the various nonlinear contributions is preferred.
@@
These general prescriptions have interpreted measurements of $\rm H$}i and galactic distributions to constrain cosmological parameters. Of the two tracers, galaxies are the traditional probe of large-scale structure (e.g., S. Cole et al., 2005), but $\rm H$}i has arisen as a promising candidate due to its ability to cheaply and quickly observe large volumes of space via 21cm intensity mapping (S. Bharadwaj et al., 2001; R. A. Battye et al., 2004; M. Leo et al., 2020). However, 21cm intensity mapping experiments still struggle with low signal-to-noise ratios, arising from galactic foregrounds and systematics (T. D. Matteo et al., 2002), making the interpretation of $\rm H$}i auto power spectra challenging (S. Paul et al., 2023, hereafter P23). These noise sources are absent in galaxy surveys, so the noise can be reduced by cross-correlating $\rm H$}i and galaxies. These cross-correlations have constrained the cosmological abundance of $\rm H$}i ($\Omega_{\mathrm{HI}}$) at low redshifts ($z\leq 1$, T.-C. Chang et al., 2010; K. W. Masui et al., 2013; E. R. Switzer et al., 2013; C. J. Anderson et al., 2018). In particular, L. Wolz et al. (2022, hereafter W22), S. Cunnington et al. (2022, C22), CHIME Collaboration et al. (2023, CHIME23), and I. P. Carucci et al. (2024, C24) have produced $\Omega_{\mathrm{HI}}$ constraints competitive to other $\rm H$}i observation techniques. As measurement and data processing techniques improve (S. Cunnington et al., 2019; I. P. Carucci et al., 2020; J. Podczerwinski & P. T. Timbie, 2024), $\rm H$}i-galaxy cross-correlations will probe other cosmological quantities in the near future (MeerKLASS Collaboration et al., 2024).
@@
However, the resulting constraints depend on how well the adopted models capture the relationship between observed tracer and matter distributions. We can evaluate these models’ performance by comparing their predictions to known $\rm H$}i, galaxy, and matter distributions in simulations. Cosmological hydrodynamical simulations now produce realistic low-redshift $\rm H$}i and galaxy distributions in sufficiently large volumes to be useful for this purpose (Y. M. Bahé et al., 2016; D. Nelson et al., 2018; B. Diemer et al., 2019; A. R. H. Stevens et al., 2019; R. Davé et al., 2020). In particular, C. K. Osinga et al. (2024) showed that the $\rm H$}i auto- and cross-correlations from IllustrisTNG agree closely with relevant observations at $z\leq 1$. We leverage the agreement between IllustrisTNG and observations to gain insight into the performance of large-scale $\rm H$}i models in two phases. First, we report IllustrisTNG’s predictions for the various model components and assess how well their simplified forms in the general model agree with the actual values. Second, we evaluate the accuracy of popular large-scale prescriptions by comparing their predictions to the correlations computed directly from IllustrisTNG.
@@ 2 Simulation data
We use the IllustrisTNG suite of cosmological magneto-hydrodynamics simulations (D. Nelson et al., 2018; A. Pillepich et al., 2018a; V. Springel et al., 2018; J. P. Naiman et al., 2018; F. Marinacci et al., 2018; D. Nelson et al., 2019). This suite offers simulations in three different box sizes, 35 h-1 cMpc, 75 h-1 cMpc, and 205 h-1 cMpc, at varying resolutions, using the Planck Collaboration et al. (2016) cosmological parameters ($\Omega_{\rm{m}}=0.3089$, $\Omega_{\rm{b}}=0.0486$, $h=0.6774$, $\sigma_{\rm{s}}=0.8159$).
@@
The simulation is evolved using the AREPO code (V. Springel, 2010), which employs a tree-PM method for gravity calculations and a Voronoi mesh-based Godunov scheme for magneto-hydrodynamics. The IllustrisTNG simulations incorporate sub-grid models to simulate unresolved processes like star formation, gas cooling, and AGN (M. Vogelsberger et al., 2013; R. Weinberger et al., 2018). The models are calibrated against a selection of observational data to accurately represent the galaxy population at low redshifts (A. Pillepich et al., 2018b). Dark matter haloes are identified using the Friends-of-Friends (FoF) algorithm (M. Davis et al., 1985), and their internal substructures, or “subhalos”, are cataloged using the Subfind algorithm (V. Springel et al., 2001).
@@
Our main analyses focus on the largest box, 205 h-1 cMpc (TNG300), at redshifts $z=0,0.5,$ and $z=1$. We discuss the impact of resolution in Section 6 and directly compare our results to the 75 h-1 cMpc box (TNG100) in Appendix A. To limit the impact of poorly resolved galaxies, we restrict our galaxy population to those containing at least 200 stellar particles, corresponding to a minimum stellar mass of $\approx 10^{9}M_{\odot}$. We limit our analysis to redshifts $z=0,0.5,$ and $1$ where the color distribution in IllustrisTNG matches observations. The agreement worsens at earlier redshifts because IllustrisTNG lacks red, star-forming, dusty galaxies (M. Donnari et al., 2019; A. Gebek et al., 2023, 2025).
@@
We separate our galaxy sample into blue and red populations to mimic how galaxies are selected in optical surveys, as they detect galaxy samples visible in certain optical bands such as blue emission-line galaxies (ELGs) and luminous red galaxies (LRGs) (S. Alam et al., 2021). We approximate these populations by categorizing galaxies as red or blue with their rest-frame grg\!-\!r color values (C. Stoughton et al., 2002): galaxies with $g\!-\!r\geq 0.6$ are blue, with all others being red. This threshold changes with time to reflect the overall color evolution of the galaxy population. Specifically, grg\!-\!r thresholds are set at 0.6, 0.55, and 0.5 for redshifts 0, 0.5, and 1, respectively. While testing various color definitions with and without dust corrections (D. Nelson et al., 2018), C. K. Osinga et al. (2024) found that these adjustments only substantially impact small-scale clustering at $k\gtrsim 1$ h cMpc-1 (see their appendix A). A more exact method of matching IllustrisTNG galaxies to ELGs and LRGs will only somewhat impact the quantitative results from Section 4 (e.g., Tables 1-3). The scale-dependent behavior in Section 4 and overall model performance presented in Section 5 are insensitive to small changes in the definitions of blue and red galaxies.
@@
The final tracer, $\rm H$}i, is not explicitly modeled in IllustrisTNG, so we must turn to post-processing to separate atomic and molecular hydrogen (B. Diemer et al., 2018). These post-processed distributions accurately reproduce a number of observed $\rm H$}i relationships (A. R. H. Stevens et al., 2019). For example, simulated $\rm H$}i rotation velocities agree with observations on average (J. S. Goddy et al., 2023) despite some increased asymmetry (A. B. Watts et al., 2020). $\rm H$}i surface densities also match, although $\rm H$}i-rich galaxies in IllustrisTNG appear to have suppressed $\rm H$}i abundance in their centers (B. Diemer et al., 2019). We expect these small differences to be negligible in our results. We have tested a wide range of post-processing models, but find negligible difference between the different treatments (an offset of $\lesssim 2\%$ in amplitude, see C. K. Osinga et al., 2024), consequently we restrict our analysis to the single model from F. Villaescusa-Navarro et al. (2018), hereafter VN18.
@@ 3 Models of large-scale $\rm H$}i distributions
$\langle\tilde{\delta}_{i}(\boldsymbol{k})\tilde{\delta}_{j}(\boldsymbol{k}^{\prime})\rangle=(2\pi)^{3}P_{i\times j}(k)\delta_{D}^{3}(\boldsymbol{k}+\boldsymbol{k}^{\prime})\,.$ (1)
@@ 3.1 Nonlinearities
At large scales, prescriptions are relatively simple since the growth of the small large-scale perturbations ($\delta\ll 1$) can be described linearly (for derivation, see S. Dodelson & F. Schmidt, 2020). However, this growth becomes increasingly nonlinear toward small scales until gravitational instabilities form collapsed structures. On these scales, we rely on numerical methods to predict the matter power spectrum.
The transition between the linear and nonlinear regime can be found by comparing their power spectra in Fig. 1. We use camb (A. Lewis et al., 2000; A. Lewis & A. Challinor, 2011) and halofit (R. Takahashi et al., 2012; R. E. Smith & R. E. Angulo, 2019) to determine the linear ($P_{\mathrm{L}}$) and nonlinear ($P_{\mathrm{NL}}$) power spectra, respectively. The two diverge at $k\sim 0.1$ h cMpc-1 at $z=1$, which agrees with other works studying the linear-nonlinear transition (e.g., R. E. Smith et al., 2003). Roughly at the same scales, the wiggles from baryonic acoustic oscillations (D. J. Eisenstein et al., 2005) imprint on both the linear and nonlinear power spectra.
We also compare the nonlinear power spectra calculated from TNG300 and halofit. At large scales, the two roughly agree, although the small number of modes at the largest scales that TNG300 probes causes statistical variations that fall above and below the halofit result (we address this in Section 4.1). At smaller scales, halofit and TNG300 diverge, with TNG300 having slightly more power, although the shapes match closely. Differences are expected since the halofit results are tuned to dark matter-only N-body simulations and the TNG300 power spectra include baryonic effects and are subject to cosmic variance (C. J. Miller et al., 2002; R. S. Somerville et al., 2004; M. P. van Daalen et al., 2011; Y. Li et al., 2014; V. Springel et al., 2018). For demonstrative purposes, we will use the halofit matter power spectrum for this section but strictly use the TNG300 result for the remainder of the paper.
@@ 3.2 From matter to tracers
The impact of the $\rm H$}i and galaxy biases is shown with the purple line in Fig. 1, which corresponds mathematically to $b_{\mathrm{HI}}(k)b_{\mathrm{Gal}}(k)P_{\mathrm{NL}}(k)$. The ratios in the bottom panel of Fig. 1 show that both biases are constant on large scales until $k\sim 0.1$ h cMpc-1 where they increase. The galaxy and $\rm H$}i biases inherit this shape from the halo bias — halos tend to occupy the most nonlinear regions of space and thus are increasingly more clustered than matter on small scales, such that the halo bias increases (D. Jeong & E. Komatsu, 2009; T. Nishimichi, 2012; A. J. Nishizawa et al., 2013; A. Paranjape et al., 2013). Nearly all galaxies and $\rm H$}i inhabit halos at $z\leq 1$ (S. D. M. White & M. J. Rees, 1978, VN18) such that their biases mirror the scale-dependence of the halo bias. The $\rm H$}i and galaxy biases used in Fig. 1 are estimated from TNG300 (discussed more in Section 4) — for scales beyond what TNG300 probes, we assume that the biases are constant and extrapolate the large-scale values from TNG300.
@@
Similarly to the bias, the correlation coefficient is expected to be constant on large scales and approach zero toward small scales as stochasticity becomes more significant (I. P. Carucci et al., 2017; C. K. Osinga et al., 2024). These trends are shown with the maroon line in Fig. 1, which represents the expression $b_{\mathrm{HI}}(k)b_{\mathrm{Gal}}(k)r_{\mathrm{HI-Gal}}(k)P_{\mathrm{NL}}(k)$ and completely describes the tracer-tracer and tracer-matter relationships. We use the $\rm H$}i-galaxy correlation coefficient from TNG300, assuming that $r_{\mathrm{HI-Gal}}=1$ at scales larger than what TNG300 probes. The correlation coefficient suppresses the power spectrum at $k\gtrsim 0.2$ h cMpc-1, with this effect strengthening toward small scales. Overall, the contribution of the correlation coefficient is fairly small relative to the other components, reducing the power spectrum by a factor of 2 at the smallest scales shown in Fig. 1.